A bullet of mass m = 5.00 g is fired into a block of mass M = 500 g initially at rest at the edge of a frictionless table of height h = 1.20 m. The bullet remains in the block, and after impact the block lands a distance d = 0.80 m from the bottom of the table. Determine the initial speed of the bullet.
Answer #1

here,

height of table , h = 1.2 m

let the time taken to hit the floor be t

h = 0 + 0.5 * g * t^2

1.2 = 0 + 0.5 * 9.81 * t^2

solving for t

t = 0.49 s

let the final speed be v

v * t = d

v * 0.49 = 0.8

v = 1.62 m/s

m = 5 g , M = 500 g

let the initial speed be u

using conservation of momentum

m * u = ( m + M) * v

5 * u = ( 500 + 5) * 1.62

solving for u

u = 163.4 m/s

the initial speed of bullet is 163.4 m/s